Pearl Products Limited of Shenzhen, China, manufactures and distributes toys throughout South East Asia. Three cubic centimeters (cc) of solvent H300 are required to manufacture each unit of Supermix, one of the company’s products. The company now is planning raw materials needs for the third quarter, the quarter in which peak sales of Supermix occur. To keep production and sales moving smoothly, the company has the following inventory requirements: The finished goods inventory on hand at the end of each month must equal 2,000 units of Supermix plus 20% of the next month’s sales. The finished goods inventory on June 30 is budgeted to be 15,400 units. The raw materials inventory on hand at the end of each month must equal one-half of the following month’s production needs for raw materials. The raw materials inventory on June 30 is budgeted to be 102,000 cc of solvent H300. The company maintains no work in process inventories. A monthly sales budget for Supermix for the third and fourth quarters of the year follows. Budgeted Unit Sales July 67,000 August 72,000 September 82,000 October 62,000 November 52,000 December 42,000 Required: 1. Prepare a production budget for Supermix for the months July, August, September, and October. 3. Prepare a direct materials budget showing the quantity of solvent H300 to be purchased for July, August, and September, and for the quarter in total.
Answer #1
Required 1 :
Pearl Products Limited
Production Budget
July August September October
Budgeted unit sales 67000 72000 82000 62000
(+) Desired units of ending finished goods inventory ( 2000 + 20% of next month's sales ) 16400 18400 14400 12400
Total needs 83400 90400 96400 74400
(-) Units of Beginning finished goods inventory 15400 16400 18400 14400
Required production in units 68000 74000 78000 60000
Note : October's ending inventory = 2000 + ( 20% * November unit sales ) = 2000 + ( 20% * 52000 ) 12400
Required 2 :
Preal Products Limited
Direct materials budget
July August September Third Quarter
Required production in units 68000 74000 78000 220000
Units of raw materials needed per unit of finished goods 3 cc 3 cc 3 cc 3 cc
Units of raw materials needed to meet production 204000 222000 234000 660000
(+) Desired units of ending raw materials inventory ( One-half of following month's production needs ) 111000 117000 90000 90000
Total units of raw materials needed 315000 339000 324000 750000
(-) Units of Beginning raw materials inventory 102000 111000 117000 102000
Units of raw materials to be purchased 213000 228000 207000 648000
Note : September's ending inventory = 1/2 * October month's production needs = 1/2 * ( 60000 * 3 ) 90000